OpenStudy (anonymous):

Can someone help me with this? √45n^5

OpenStudy (solomonzelman):

\(\large\sqrt{45n^5}\) like this?

OpenStudy (anonymous):

Yeah like that.

OpenStudy (solomonzelman):

\(\color{black}{ \displaystyle \sqrt{45n^5}=\sqrt{9\cdot 5 \cdot n^4\cdot n} =\sqrt{9\cdot n^4}~\cdot ~\sqrt{5\cdot n}=\sqrt{9} \cdot \sqrt{ n^4}~\cdot ~\sqrt{5\cdot n} }\)

OpenStudy (solomonzelman):

that is how i brake them down.

OpenStudy (solomonzelman):

what is √9, can you tell me?

OpenStudy (anonymous):

3?

OpenStudy (solomonzelman):

yes

OpenStudy (anonymous):

How did we lose an n from n^5?

OpenStudy (solomonzelman):

we didn't, all I did is used the fact that \(n^5=n^{4+1}=n^4\cdot n^1=n^4\cdot n\)

OpenStudy (solomonzelman):

and of course, I rearranged them putting the terms that can be simplified in the beginning and then the terms that can't be simplified

OpenStudy (solomonzelman):

\(\sqrt{n^4}\) is same as \(\sqrt[2]{n^4}\) and that is =\(n^{4/2}={\rm to~what?}\)

OpenStudy (anonymous):

Oh alright..

OpenStudy (solomonzelman):

ok, did you get up to finding the \(\sqrt{n^4}\) ?

OpenStudy (anonymous):

Nope don't understand what's going on this time.

OpenStudy (solomonzelman):

\(\sqrt{n^4}=\sqrt[2]{n^4}\) this is because whenever the root is not specificed, it is 2 (i.e. SQUARE root)

OpenStudy (solomonzelman):

now, applying the rule: \(\sqrt[c]{a^b}=a^{b/c}\) we get that \(\sqrt[2]{n^4}=n^{4/2}=n^{2/1}=n^2\)

OpenStudy (anonymous):

Okay, didn't know that.

OpenStudy (solomonzelman):

We had: \(\color{black}{ \displaystyle \sqrt{9} \cdot \sqrt{ n^4}~\cdot ~\sqrt{5\cdot n} }\) and now we had: \(\color{black}{ \displaystyle3 \cdot n^2\cdot ~\sqrt{5\cdot n} }\)

OpenStudy (solomonzelman):

oh, now you know.... :0

OpenStudy (anonymous):

Thanks.

OpenStudy (solomonzelman):

\(\color{black}{ \displaystyle3 \cdot n^2\cdot ~\sqrt{5\cdot n} }\) is just same as \(\color{black}{ \displaystyle3 n^2~\sqrt{5 n} }\) \(\bf ... \)yw